OpenStudy (day_2010):

How do you solve this: Ellen wishes to mix candy worth $1.64 per pound with candy worth $3.36 per pounds to form 30 pounds of a mixture worth $2.33 per pound. How many pounds of the more expensive candy should she use? help please! :)

OpenStudy (anonymous):

start by determining the total cost of the 30 pounds of candy, what do you get?

OpenStudy (day_2010):

1.25

OpenStudy (anonymous):

Let cheap candy quantity be C. Let expensive candy quanity be E. C + E = 30 (1.64C + 3.36E)/30 = 2.33 Solve for C and E.

OpenStudy (anonymous):

C + E = 30 1.64C + 3.36E = 69.9 1.64(30-E) + 3.36E = 69.9 49.2 - 1.64E + 3.36E = 69.9 3.36E - 1.64E = 69.9 - 49.2 1.72E = 20.7 E = 20.7/1.72 = ?

OpenStudy (anonymous):

12 Pounds.

OpenStudy (day_2010):

Oh okay... thanks for simplifying this for me... I hate word problems and this one gave me trouble. Thank You So Much!

OpenStudy (anonymous):

no, that is the one i tried and it didn't work. sorry, i will delete

OpenStudy (anonymous):

No. That won't do it.

OpenStudy (anonymous):

LOL -- the next line in my notes show the same you provided.